OpenStudy (anonymous):

I7x-8I less than 4. express using interval notation

OpenStudy (anonymous):

|a|<b can be written as -b < a < b. So in your case, |7x - 8| < 4 is written as -4 < 7x - 8 < 4 Add an 8 throughout, and then divide by 7. To put in interval notation, remember that a < x < b is written as (a,b)

OpenStudy (anonymous):

how should that look? i dont know if i did it right

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I'll do the first step to get you started :) -4 < 7x - 8 < 4 4 < 7x < 12 Now what would the next line look like when you divide through by 7?

OpenStudy (anonymous):

4/7 less than x less than 12/7?

OpenStudy (anonymous):

right! now for the interval notation... if you have a < x < b, your interval notation looks like (a,b) meaning all values between a and b. you have 4/7 < x < 12/7, so the interval notation would be...

OpenStudy (anonymous):

(4/7,12/7)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

perfect! :)

OpenStudy (anonymous):

could u help with my other ? too

OpenStudy (anonymous):

sorry...had to take a phone call. what's the question?

OpenStudy (anonymous):

(4/x)-(40/x+6)+3=0. find all real solutions

OpenStudy (anonymous):

just like adding and subtracting regular fractions, you need a common denominator. the denominators of the three fractions you have are x, x+6, and 1

OpenStudy (anonymous):

one way to always find a common denominator is to multiply them together. you can do this with regular fractions too, but it often gives you a bigger number than you need. in this case, however, it's all we can do. so, your common denominator is x(x+6). don't bother multiplying it out...it's nicer in this form for now

OpenStudy (anonymous):

your first fraction's denominator is missing the x+6, so we multiply top and bottom by it. likewise, the second fraction is missing the x. the thrid fraction is missing both x and x+6, so we multiply top and bottom by the whole thing. this gives us: 4(x+6) 40x 3x(x+6) --------- - ------ + -------- = 0 x(x+6) x(x+6) x(x+6)

OpenStudy (anonymous):

Just like with regular fractions, once you have a common denominator, you can add the numerators and put that over the common denominator.

OpenStudy (anonymous):

you can multiply out the individual numerators here, but leave the denominator as is for now. 4x + 24 - 40x + 3x^2 + 18x -------------------------- = 0 x(x+6)

OpenStudy (anonymous):

A fraction will be 0 only when the numerator is 0, so to solve this, it must be that \[4x+24-40x+3x^{2}+18=0\]simplify the equation\[3x^{2}-36x+42=0\]

OpenStudy (anonymous):

since every term has a 3 in it, divide through by 3.\[x^{2}-12x+14=0\]

OpenStudy (anonymous):

have you learned how to solve quadratic equations? try to see what you get.